kate walks 5 miles in 2 hours at a steady pace.How far can she walk in 1 hour and 15 minutes?
{please also explain how you got the answer)

Answer by stanbon(75887)

kate walks 5 miles in 2 hours at a steady pace.How far can she walk in 1 hour and 15 minutes?
{please also explain how you got the answer)
-----
Use a proportion::
dist/time = dist/time
---
d/75min = 5/120min
--------------------------
d = 75*5/120
---
distance = 3.125 = 3 1/8 mile
------------------
Cheers,
Stan H.